OpenStudy (anonymous):

Bonita evaluated 5/6 × 1 2/3 and got an answer of 7/3. How can you tell that her answer is wrong?

OpenStudy (anonymous):

@Skydog07 may you help me? :))

OpenStudy (immanuelv):

you can tell she got the wrong answer because a) she didn't use a calculator b) she didn't do her calculation right

OpenStudy (immanuelv):

Bonita should've multiplied \[\frac{5 }{ 6 } \times \frac{ 5 }{ 3 } = \frac{25}{18}\]

OpenStudy (dtan5457):

Well it's wrong simply cause it's wrong to be honest. If she had put them back to improper fractions, 5/6 times 5/4 is 25/18. 18 divided by 6 is 3, but 25 divided by 6 is obviously not 7.

OpenStudy (anonymous):

1 2/3 = 1 4/6. 7/3 would equal 2 1/3 or 2 2/6. 5/6 times 5/3 would equal 25/18, which is 1 7/18 which not is equal to 21/18, which is equal to 7/3. sorry if its confusing